Daily Enterprises is purchasing a $10.57 million machine. It will cost $59,244.00 to transport and install the machine. The machine has a depreciable life of five years using the straight-line depreciation and will have no salvage value. The machine will generate incremental revenues of $4.31 million per year along with incremental costs of $1.20 million per year. Daily’s marginal tax rate is 34.00%.

The cost of capital for the firm is 10.00%.

(answer in dollars..so convert millions to dollars)

A) What is the year 0 cash flow for the project?

B) What is the yearly cash flow from the project?

C) The project will run for 5 years. What is the NPV of the project at the current cost of capital?
Answer #1

Answer 1  
Year 0 Cash flow  
  
Cost of machine   -$10,570,000
transport and Installation cost   -59244
___________________________  
Total Cash flow at year 0   -$10,629,244
___________________________  

Cash flow from Year 1 to 5  
   Year 1
Incremental revenue   4310000.00
less: cost   -1200000.00
less: Depreciation   -2125848.80
(10629244/5)  
___________________________  
Operating profit   984151.20
less tax @ 34%   -334611.41
___________________________  
Profit after tax   649539.79
Add: Depreciation   2125848.80
___________________________  
Free cash flow   2775388.59
  
  
(2) Depreciation is non cash expenses. It is only deducted for tax calculation. Again it will be added as it is not in cash.

NpV = sum of all free cash flows          
PVF formula = 1/(1+i)^n          
for year 1 PVF =1/(1+10%)^1=   0.9090909091      
for year 2 PVF =1/(1+10%)^2=   0.826446281   and So on.  
          
Calcultion of NPV          
Year   Free cash flow   PVF @10%   PV = Cash flow * PVF
0   -$10,629,244   1.0000   -10,629,244.00
1   2775388.59   0.9091   2,523,080.54
2   2775388.59   0.8264   2,293,709.58
3   2775388.59   0.7513   2,085,190.53
4   2775388.59   0.6830   1,895,627.75
5   2775388.59   0.6209   1,723,297.96
          
NpV = sum of all free cash flows=           -108,337.65
          
So NPV of project is    -108,337.65      
      

А Answer 1 Year O Cash flow 3 4 - Cost of machine transport and Ins -$10,570,000 -59244 Total Cash flow a $10,629,244 10 11 7А 45 46 47 48 NPV = sum of all free cash flows PVF formula = 1/(1+i)^n for year 1 PVF = 0.9090909091 for year 2 PVF = 0.82644